# SpikingBrain: Spiking Brain-inspired Large Models

### v2

@@ Abstract
Mainstream Transformer-based large language models (LLMs) face significant efficiency bottlenecks: training computation scales quadratically with sequence length, and inference memory grows linearly. These constraints limit their ability to process long sequences effectively. In addition, building large models on non-NVIDIA computing platforms poses major challenges in achieving stable and efficient training and deployment. To address these issues, we introduce SpikingBrain, a new family of brain-inspired models designed for efficient long-context training and inference. SpikingBrain leverages the MetaX [^1] https://www.metax-tech.com/en GPU cluster and focuses on three core aspects: i) Model Architecture: linear and hybrid-linear attention architectures with adaptive spiking neurons; ii) Algorithmic Optimizations: an efficient, conversion-based training pipeline compatible with existing LLMs, along with a dedicated spike coding framework; iii) System Engineering: customized training frameworks, operator libraries, and parallelism strategies tailored to the MetaX hardware.
Using these techniques, we develop two models: SpikingBrain-7B, a linear LLM, and SpikingBrain-76B, a hybrid-linear MoE LLM. These models demonstrate the feasibility of large-scale LLM development on non-NVIDIA platforms. SpikingBrain achieves performance comparable to open-source Transformer baselines while using exceptionally low data resources (continual pre-training of $\sim$150B tokens). Our models also significantly improve long-sequence training efficiency and deliver inference with (partially) constant memory and event-driven spiking behavior. For example, SpikingBrain-7B achieves more than 100× speedup in Time to First Token (TTFT) for 4M-token sequences. Our training framework supports weeks of stable large-scale training on hundreds of MetaX C550 GPUs, with the 7B model reaching a Model FLOPs Utilization (MFU) of 23.4%. In addition, the proposed spiking scheme achieves 69.15% sparsity, enabling low-power operation. Overall, this work demonstrates the potential of brain-inspired mechanisms to drive the next generation of efficient and scalable large model design. [^2] The code for this project is publicly available at https://github.com/BICLab/SpikingBrain-7B.
@@ 1 Introduction
Recent advances in large language models (LLMs) built on the Transformer architecture (vaswani2017attention, ) have been driven by the scaling law (kaplan2020scaling, ), which suggests that performance improves with larger model sizes and more data (OpenAI2025GPT5, ; GoogleDeepMind2024GeminiPro, ; Anthropic2025ClaudeOpus41, ; liu2024deepseek, ). However, this scale-driven approach comes with significant challenges: extremely high training costs, substantial energy consumption, and complex deployment pipelines. Therefore, achieving high performance and energy efficiency under limited resources has become a critical research goal. To address this, our work draws inspiration from brain mechanisms. We explore novel architectures, training paradigms, and spike coding schemes to develop efficient, brain-inspired LLMs that move beyond the traditional Transformer framework.
A further objective is to validate the training and deployment of such models on non-NVIDIA computing clusters. We use an open-source Transformer checkpoint (Qwen2.5-7B-base (yang2024qwen2, ) as an example) together with our efficient development framework to train and evaluate two models on the MetaX GPU cluster. The models, SpikingBrain-7B and SpikingBrain-76B-A12B, undergo end-to-end validation, including continual pre-training (CPT), long-context extension (up to 128k tokens), and supervised fine-tuning (SFT). We also adapt the vLLM inference framework to demonstrate deployment feasibility on MetaX hardware.
@@
  • Hybrid Linear Architectures: Moving away from quadratic self-attention, we design hybrid models combining linear, local, and standard attention modules. We explore two hybrid strategies: inter-layer sequential (SpikingBrain-7B) and intra-layer parallel (SpikingBrain-76B-A12B). The former achieves linear complexity and excels at long-context efficiency; the latter provides stronger language modeling capability through a more sophisticated architectural design. Notably, the modeling of linear attention (katharopoulos2020transformers, ) also resonates with neuronal dynamics in biological systems (he2024network, ).
  • Adaptive Threshold Spiking: Inspired by event-driven biological neurons, we propose a spiking scheme that converts activations into integer spike counts and expands them into sparse spike trains. This enables addition-based, event-driven computation. Several encoding formats are supported, including binary {0,1}, ternary {-1,0,1}, and bitwise spike coding. These sparse, event-driven representations provide the basis for our low-power inference design and may also inspire the development of next-generation neuromorphic hardware (roy2019towards, ; schuman2022opportunities, ; frenkel2023bottom, ).
  • Efficient Model Conversion: i) Attention modules: Using a unified attention map analysis, we convert quadratic attention modules into sparse sliding-window and low-rank linear attention, by remapping the weights of existing Transformer models (kasai2021finetuning, ). This reduces training and inference costs, enabling efficient long-context handling with less than 2% of the compute needed for training from scratch. ii) FFN modules: For SpikingBrain-76B, we utilize an MoE upcycling technique (he2024upcycling, ) that replicates dense FFN weights to create sparse experts, increasing capacity with minimal compute and memory overhead.
@@
Our design philosophy holds that linear attention modules exhibit modeling characteristics highly analogous to human memory mechanisms, particularly through their use of compressed, continuously updated memory states (gu2025tradeoffs, ). From a biological perspective, linear attention can also be interpreted as a simplified form of dendritic dynamics. In addition, the Mixture-of-Experts (MoE) component reflects a principle of modular specialization, akin to the distributed and specialized processing observed in biological neural networks (o2021and, ). By combining network-level sparsity (via MoE) with neuron-level spiking sparsity, our approach offers a robust multi-scale efficiency strategy. Taken together, these results point toward a promising direction for developing more efficient and biologically plausible large-model architectures that extend beyond standard Transformer-based LLMs through brain-inspired mechanisms.
In summary, this work presents efficient brain-inspired LLM development on MetaX GPUs, producing two models: SpikingBrain-7B and SpikingBrain-76B MoE. They achieve (near-)linear complexity, high training stability, and competitive performance with only $\sim 2\%$ of the pre-training data typical for mainstream LLMs. Inference shows substantial speedups, with SpikingBrain-7B reaching over 100× speedup in TTFT (Time to First Token) for 4M-token inputs. We further deploy a compressed 1B SpikingBrain model on a CPU-based mobile inference framework, achieving a 15.39× speedup at a sequence length of 256k. The proposed spiking scheme also yields $\sim 69\%$ sparsity, offering strong support for reducing power consumption. These findings highlight the potential of brain-inspired design for future neuromorphic hardware and enable efficient model scaling across diverse computing platforms (see Figure 2).
Figure 2: Compatibility of SpikingBrain models across diverse computing platforms. SpikingBrain models can be deployed on CPUs and both NVIDIA and non-NVIDIA GPUs using integer activation formats, and they also inspire the design of dedicated neuromorphic hardware leveraging event-driven sparse spike representations.
@@ 2.1.1 Attention Mechanisms
Standard softmax attention (vaswani2017attention, ) captures global, token-wise interactions across the entire sequence:
@@
Here, $\mathbf{M}$ is a causal mask where $\mathbf{M}_{ij}=1$ if $i\geq j$ and $\mathbf{M}_{ij}=$-$\infty$ if $i<j$. For a sequence of length $n$, softmax attention provides high arithmetic intensity and parallelism during training but incurs $O(n^{2})$ computation. During inference, maintaining the key-value cache adds an $O(n)$ memory footprint, becoming a bottleneck for long-context tasks.
@@
To reduce this quadratic computational cost, many linear-complexity variants have been proposed. One example is SWA (beltagy2020longformer, ; child2019generating, ; Jiang2023Mistral7, ), which limits the attention scope to a fixed local window of size $w$, capturing fine-grained local interactions:
@@
Here, $\mathbf{M^{\prime}}$ is a windowed causal mask where $\mathbf{M^{\prime}}_{ij}=1$ if $i-w+1\leq j\leq i$, otherwise $\mathbf{M^{\prime}}_{ij}=$-$\infty$. Thus, training complexity and inference memory reduce to $O(n)$ and $O(1)$ respectively, as $w$ is a fixed constant independent of $n$.
@@
Another common variant is linear attention (katharopoulos2020transformers, ; yang2024gated, ; dao2024transformers, ), which removes the softmax function to achieve linear complexity and can be reformulated as a state-based linear recurrence:
@@
Here, $\mathbf{M}$ is a causal mask where $\mathbf{M}_{ij}=1$ if $i\geq j$ and $\mathbf{M}_{ij}=0$ if $i<j$. Linear attention supports chunk-wise parallelism during training and maintains a fixed-size state in recurrent form (sun2023retentive, ; yang2024gated, ), reducing computational complexity and inference memory to $O(n)$ and $O(1)$.
@@
Each attention mechanism has distinct strengths: softmax attention excels at global retrieval, SWA is efficient for local context, and linear attention effectively compresses long-range information. Hybrid Attention aims to combine these methods to balance efficiency and accuracy. Two common paradigms are: i) Inter-layer sequential hybridization (lieber2024jamba, ; rensamba, ; de2024griffin, ), where different attention types are stacked across layers:
@@
ii) Intra-layer parallel hybridization (dong2024hymba, ; zuo2025falcon, ), where different attention modules process the same input in parallel and outputs are merged:
@@ 2.1.2 Mixture-of-Experts (MoE)
The core idea of sparse MoE (lepikhin2020gshard, ; fedus2022switch, ) is to enhance model capacity by introducing $N$ parallel expert networks into the original feed-forward network (FFN) layer, while dynamically selecting the most relevant $k$ experts for each input token $\mathbf{x}$ through a router $\mathbf{W}_{r}$:
@@
Unlike traditional dense layers, MoE does not activate all experts simultaneously. Instead, it leverages sparse activation to significantly increase parameter capacity and expressiveness while keeping computational cost nearly unchanged. This property enables MoE to combine efficiency with high performance in both training and inference. Prior studies (dai2024deepseekmoe, ; Yang2024Qwen2TR, ; liu2024deepseek, ) further suggest that introducing shared experts that are always active, and retaining a few dense FFN layers in shallow stages of the model can improve training stability and overall performance.
In practice, dense models can be efficiently expanded into MoE models using the upcycling technique (komatsuzaki2022sparse, ; he2024upcycling, ), which scales parameter capacity without sacrificing the original performance. The procedure involves: i) replicating the dense FFN weights across all experts so that the upcycled model initially matches the dense baseline; ii) rescaling expert outputs appropriately to maintain consistency with the output scale of the original model.
@@ 2.1.3 Spiking Neuron Modeling
Spiking neurons are the core components of Spiking Neural Networks (SNNs) (maass1997networks, ), and their modeling generally falls into two paradigms: simplified approximations and detailed simulations. The Hodgkin–Huxley (HH) model (hodgkin1952quantitative, ) represents the latter, accurately describing the membrane potential dynamics of biological neurons through a set of nonlinear differential equations. It focuses on simulating the synergistic effects of transmembrane ion fluxes. Although such models can faithfully replicate fine-scale bioelectrical phenomena of neurons (e.g., complex firing patterns and refractory periods), they entail high computational complexity due to the involvement of multiple high-order differential operations. Thus, HH-based models are currently mainly suitable for physiological studies on small neuron populations and cannot meet the efficiency and latency requirements of LLMs.
By contrast, simplified models such as the Leaky Integrate-and-Fire (LIF) model, are more commonly used in practice (yao2023spike, ; meta_spikeformer, ). The LIF neuron is a first-order approximation of soma dynamics incorporating the temporal dimension. For an input token $\mathbf{x}$ (extended over time, where $\mathbf{x}_{t}$ is the input at the $t$-th time step), the membrane potential $\mathbf{v}_{t}$ and spike output $\mathbf{s}_{t}$ of the neuron can be formulated as below:
$\mathbf{v}_{t+1}=\begin{cases}\lambda(1-\mathbf{s}_{t})\mathbf{v}_{t}+\mathbf{v}_{\text{reset}}\cdot\mathbf{s}_{t}+\mathbf{x}_{t+1},&\text{hard reset}\\ \lambda\mathbf{v}_{t}-V_{\text{th}}\cdot\mathbf{s}_{t}+\mathbf{x}_{t+1},&\text{soft reset}\end{cases};\quad\mathbf{s}_{t}=1\ \text{ if }\ \mathbf{v}_{t}\geq V_{\text{th}}\ \text{ else }\ 0.$ (9)
Here, the membrane potential accumulates charge in the soma, $\lambda$ represents a decay factor (mimicking the natural leakage of ion potential), where $V_{\text{th}}$ represents a fixed firing threshold, and $\mathbf{v}_{\text{reset}}$ is the reset potential, which is usually set to 0. While this model simplifies the ion-based mechanisms and retains the core logic of natural neurons, it still exhibits several limitations for large-scale models (luo2025integer, ; yao2024scaling, ; Fan2025Multisynaptic, ): i) The temporal dynamics (mainly introduced by the decay factor and reset mechanism) lead to increased training complexity and instability; ii) Even when integrated into pre-trained models, redundant complexity still persists, making it complicated to build models and simulate the original values; iii) The fixed threshold can cause suboptimal spike generation, leading to many neurons becoming silent or over-activated, posing challenges for optimization and impeding the balance between accuracy and energy efficiency.
@@
  • Simplified Temporal Computation: The decay factor is eliminated, and the soft-reset mechanism is adopted, enabling the conversion from continuous values to integer spike counts in a single step. During optimization, the temporal dimension is merged for stability and computational efficiency on GPUs. For inference, the temporal dimension can be re-expanded, and energy-efficient computation can be achieved using sparse event-driven asynchronous hardware (Speck, ).

### v3

@@ Abstract
Mainstream Transformer-based large language models (LLMs) face significant efficiency bottlenecks: training computation scales quadratically with sequence length, and inference memory grows linearly. These constraints limit their ability to process long sequences effectively. In addition, building large models on non-NVIDIA computing platforms poses major challenges in achieving stable and efficient training and deployment. To address these issues, we introduce SpikingBrain, a new family of brain-inspired models designed for efficient long-context training and inference. SpikingBrain leverages the MetaX [^1] https://www.metax-tech.com/en GPU cluster and focuses on three core aspects: i) Model Architecture: linear and hybrid-linear attention architectures with adaptive spiking neurons; ii) Algorithmic Optimizations: an efficient, conversion-based training pipeline compatible with existing LLMs, along with a dedicated spike coding framework; iii) System Engineering: customized training frameworks, operator libraries, and parallelism strategies tailored to the MetaX hardware. Using these techniques, we develop two models: SpikingBrain-7B, a linear LLM, and SpikingBrain-76B, a hybrid-linear MoE LLM. These models demonstrate the feasibility of large-scale LLM development on non-NVIDIA platforms, and our training framework supports weeks of stable training on hundreds of MetaX GPUs with Model FLOPs Utilization (MFU) at expected levels. SpikingBrain achieves performance comparable to open-source Transformer baselines while using exceptionally low data resources (continual pre-training of $\sim$150B tokens). Our models also significantly improve long-context efficiency and deliver inference with (partially) constant memory and event-driven spiking behavior. For example, SpikingBrain-7B achieves more than 100× speedup in Time to First Token (TTFT) for 4M-token sequences. Furthermore, the proposed spiking scheme achieves 69.15% sparsity, enabling low-power operation. Overall, this work demonstrates the potential of brain-inspired mechanisms to drive the next generation of efficient and scalable large model design. [^2] The code for this project is publicly available at https://github.com/BICLab/SpikingBrain-7B.
@@ 1 Introduction
Recent advances in large language models (LLMs) built on the Transformer architecture (Vaswani et al., 2017) have been driven by the scaling law (Kaplan et al., 2020), which suggests that performance improves with larger model sizes and more data (OpenAI, 2025; Google DeepMind, 2025; Anthropic, 2025). However, this scale-driven approach comes with significant challenges: extremely high training costs, substantial energy consumption, and complex deployment pipelines. Therefore, achieving high performance and energy efficiency under limited resources has become a critical research goal. To address this, our work draws inspiration from brain mechanisms. We explore novel architectures, training paradigms, and spike coding schemes to develop efficient, brain-inspired LLMs that move beyond the traditional Transformer framework.
A further objective is to validate the training and deployment of such models on non-NVIDIA computing clusters. We use an open-source Transformer checkpoint (Qwen2.5-7B-base (Yang et al., 2024b) as an example) together with our efficient development framework to train and evaluate two models on the MetaX GPU cluster. The models, SpikingBrain-7B and SpikingBrain-76B-A12B, undergo end-to-end validation, including continual pre-training (CPT), long-context extension (up to 128k tokens), and supervised fine-tuning (SFT). We also adapt the vLLM inference framework to demonstrate deployment feasibility on MetaX hardware.
@@
  • Hybrid Linear Architectures: Moving away from quadratic self-attention, we design hybrid models combining linear, local, and standard attention modules. We explore two hybrid strategies: inter-layer sequential (SpikingBrain-7B) and intra-layer parallel (SpikingBrain-76B-A12B). The former achieves linear complexity and excels at long-context efficiency; the latter provides stronger language modeling capability through a more sophisticated architectural design. Notably, the modeling of linear attention (Katharopoulos et al., 2020) also resonates with neuronal dynamics in biological systems (He et al., 2024b).
  • Adaptive Threshold Spiking: Inspired by event-driven biological neurons, we propose a spiking scheme that converts activations into integer spike counts and expands them into sparse spike trains. This enables addition-based, event-driven computation. Several encoding formats are supported, including binary $\{0,1\}$, ternary $\{-1,0,1\}$, and bitwise spike coding. These sparse, event-driven representations provide the basis for our low-power inference design and may also inspire the development of next-generation neuromorphic hardware (Roy et al., 2019; Schuman et al., 2022; Frenkel et al., 2023).
  • Efficient Model Conversion: i) Attention modules: Using a unified attention map analysis, we convert quadratic attention modules into sparse sliding-window and low-rank linear attention, by remapping the weights of existing Transformer models (Kasai et al., 2021). ii) FFN modules: For SpikingBrain-76B, we utilize an MoE upcycling technique (He et al., 2024a) that replicates dense FFN weights to create sparse experts, increasing capacity with minimal compute and memory overhead. These techniques, together with the dedicated spike coding framework, reduce continual training and inference costs, enabling efficient long-context handling with less than 2% of the compute required for training from scratch.
@@
Our design philosophy holds that linear attention modules exhibit modeling characteristics highly analogous to human memory mechanisms, particularly through their use of compressed, continuously updated memory states (Gu, 2025). From a biological perspective, linear attention can also be interpreted as a simplified form of dendritic dynamics. In addition, the Mixture-of-Experts (MoE) component reflects a principle of modular specialization, akin to the distributed and specialized processing observed in biological neural networks (O’Doherty et al., 2021). By combining network-level sparsity (via MoE) with neuron-level spiking sparsity, our approach offers a robust multi-scale efficiency strategy. Taken together, these results point toward a promising direction for developing more efficient and biologically plausible large-model architectures that extend beyond standard Transformer-based LLMs through brain-inspired mechanisms.
@@
Figure 2: Compatibility of SpikingBrain models across diverse computing platforms. SpikingBrain models can be deployed on CPUs and both NVIDIA and non-NVIDIA GPUs using integer activation formats, also inspiring the design of neuromorphic hardware leveraging event-driven sparse spike representations.
In summary, this work presents efficient brain-inspired LLM development on MetaX GPUs, producing two models: SpikingBrain-7B and SpikingBrain-76B MoE. They achieve (near-)linear complexity, high training stability, and competitive performance with only $\sim 2\%$ of the training data typical for mainstream LLMs. Inference shows substantial speedups, with SpikingBrain-7B reaching over 100× speedup in TTFT (Time to First Token) for 4M-token inputs. We further deploy a compressed 1B SpikingBrain model on a CPU-based mobile inference framework, achieving a 15.39× speedup at a sequence length of 256k. The proposed spiking scheme also yields $\sim 69\%$ sparsity, offering strong support for reducing power consumption. These findings highlight the potential of brain-inspired design for future neuromorphic hardware and enable efficient model scaling across diverse computing platforms (see Figure 2).
@@ 2.1.1 Attention Mechanisms
Standard softmax attention (Vaswani et al., 2017) captures global, token-wise interactions across the entire sequence:
@@
Here, $\mathbf{M}$ is a causal mask where $\mathbf{M}_{ij}=1$ if $i\geq j$ and $\mathbf{M}_{ij}=$-$\infty$ if $i<j$. For a sequence of length $n$, softmax attention provides high arithmetic intensity and parallelism during training but incurs $\mathcal{O}(n^{2})$ computation. During inference, maintaining the key-value cache adds an $\mathcal{O}(n)$ memory footprint, becoming a bottleneck for long-context tasks.
@@
To reduce this quadratic computational cost, many linear-complexity variants have been proposed. One example is SWA (Beltagy et al., 2020; Child et al., 2019; Jiang et al., 2023), which limits the attention scope to a fixed local window of size $w$, capturing fine-grained local interactions:
@@
Here, $\mathbf{M^{\prime}}$ is a windowed causal mask where $\mathbf{M^{\prime}}_{ij}=1$ if $i-w+1\leq j\leq i$, otherwise $\mathbf{M^{\prime}}_{ij}=$-$\infty$. Thus, training complexity and inference memory reduce to $\mathcal{O}(n)$ and $\mathcal{O}(1)$ respectively, as $w$ is a fixed constant independent of $n$.
@@
Another common variant is linear attention (Katharopoulos et al., 2020; Yang et al., 2024c; Dao & Gu, 2024), which removes the softmax function to achieve linear complexity and can be reformulated as a state-based linear recurrence:
@@
Here, $\mathbf{M}$ is a causal mask where $\mathbf{M}_{ij}=1$ if $i\geq j$ and $\mathbf{M}_{ij}=0$ if $i<j$. Linear attention supports chunk-wise parallelism during training and maintains a fixed-size state in recurrent form (Sun et al., 2023; Yang et al., 2024c), reducing computational complexity and inference memory to $\mathcal{O}(n)$ and $\mathcal{O}(1)$.
@@
Each attention mechanism has distinct strengths: softmax attention excels at global retrieval, SWA is efficient for local context, and linear attention effectively compresses long-range information. Hybrid Attention aims to combine these methods to balance efficiency and accuracy. Two common paradigms are: i) Inter-layer sequential hybridization (Lieber et al., 2024; Ren et al., 2024; De et al., 2024), where different attention types are stacked across layers:
@@
ii) Intra-layer parallel hybridization (Dong et al., 2024; Zuo et al., 2025), where different attention modules process the same input in parallel and outputs are merged:
@@ 2.1.2 Mixture-of-Experts (MoE)
The core idea of sparse MoE (Lepikhin et al., 2020; Fedus et al., 2022) is to enhance model capacity by introducing $N$ parallel expert networks into the original feed-forward network (FFN) layer, while dynamically selecting the most relevant $k$ experts for each input token $\mathbf{x}$ through a router $\mathbf{W}_{r}$:
@@
Unlike traditional dense layers, MoE does not activate all experts simultaneously. Instead, it leverages sparse activation to significantly increase parameter capacity and expressiveness while keeping computational cost nearly unchanged. This property enables MoE to combine efficiency with high performance in both training and inference. Prior studies (Dai et al., 2024; Yang et al., 2024a; Liu et al., 2024) further suggest that introducing shared experts that are always active, and retaining a few dense FFN layers in shallow stages of the model can improve training stability and overall performance.
In practice, dense models can be efficiently expanded into MoE models using the upcycling technique (Komatsuzaki et al., 2022; He et al., 2024a), which scales parameter capacity without sacrificing the original performance. The procedure involves: i) replicating the dense FFN weights across all experts so that the upcycled model initially matches the dense baseline; ii) rescaling expert outputs appropriately to maintain consistency with the output scale of the original model.
@@ 2.1.3 Spiking Neuron Modeling
Spiking neurons are the core components of Spiking Neural Networks (SNNs) (Maass, 1997), and their modeling generally falls into two paradigms: simplified approximations and detailed simulations. The Hodgkin–Huxley (HH) model (Hodgkin & Huxley, 1952) represents the latter, accurately describing the membrane potential dynamics of biological neurons through a set of nonlinear differential equations. It focuses on simulating the synergistic effects of transmembrane ion fluxes. Although such models can faithfully replicate fine-scale bioelectrical phenomena of neurons (e.g., complex firing patterns and refractory periods), they entail high computational complexity due to the involvement of multiple high-order differential operations. Thus, HH-based models are currently mainly suitable for physiological studies on small neuron populations and cannot meet the efficiency and latency requirements of LLMs.
By contrast, simplified models such as the Leaky Integrate-and-Fire (LIF) model, are more commonly used in practice (Yao et al., 2023; 2024a). The LIF neuron is a first-order approximation of soma dynamics incorporating the temporal dimension. For an input token $\mathbf{x}$ (extended over time, where $\mathbf{x}_{t}$ is the input at the $t$-th time step), the membrane potential $\mathbf{v}_{t}$ and spike output $\mathbf{s}_{t}$ of the neuron can be formulated as below:
$\mathbf{v}_{t+1}=\begin{cases}\lambda(1-\mathbf{s}_{t})\mathbf{v}_{t}+\mathbf{v}_{\text{reset}}\cdot\mathbf{s}_{t}+\mathbf{x}_{t+1}&\text{hard reset}\\ \lambda\mathbf{v}_{t}-V_{\text{th}}\cdot\mathbf{s}_{t}+\mathbf{x}_{t+1}&\text{soft reset}\end{cases};\qquad\mathbf{s}_{t}=1\ \text{ if }\ \mathbf{v}_{t}\geq V_{\text{th}}\ \text{ else }\ 0.$ (9)
Here, the membrane potential accumulates charge in the soma, $\lambda$ represents a decay factor (mimicking the natural leakage of ion potential), where $V_{\text{th}}$ represents a fixed firing threshold, and $\mathbf{v}_{\text{reset}}$ is the reset potential, which is usually set to 0. While this model simplifies the ion-based mechanisms and retains the core logic of natural neurons, it still exhibits several limitations for large-scale models (Luo et al., 2025; Yao et al., 2025; Fan et al., 2025): i) The temporal dynamics (mainly introduced by the decay factor and reset mechanism) lead to increased training complexity and instability; ii) Even when integrated into pre-trained models, redundant complexity still persists, making it complicated to build models and simulate the original values; iii) The fixed threshold can cause suboptimal spike generation, leading to many neurons becoming silent or over-activated, posing challenges for optimization and impeding the balance between accuracy and energy efficiency.
@@
  • Simplified Temporal Computation: The decay factor is eliminated, and the soft-reset mechanism is adopted, enabling the conversion from continuous values to integer spike counts in a single step. During optimization, the temporal dimension is merged for stability and computational efficiency on GPUs. For inference, the temporal dimension can be re-expanded, and energy-efficient computation can be achieved using sparse event-driven asynchronous hardware (Yao et al., 2024b).